OpenStudy (anonymous):

What're the best factors when square rooting 216?

OpenStudy (campbell_st):

216 = 36 x 6 so maybe this helps

OpenStudy (anonymous):

14.6969384567 <-- This is the exact answer.. but not sure if THAT helps.. haha..

OpenStudy (anonymous):

So, it'd be 6 square root of 6?

OpenStudy (campbell_st):

\[\sqrt{216} = \sqrt{36 \times 6} = \sqrt{36} \times \sqrt{6} \]

OpenStudy (anonymous):

He gave a better explanation. =)

OpenStudy (campbell_st):

thats correct..